# Climbing the Ladder of Interpretability with Counterfactual Concept Bottleneck Models

### v1

@@ 3.3 Climbing the “Ladder of Interpretability”
      $z^{\prime}\sim\mathcal{N}(\phi_{\mu^{\prime}}(\alpha),\phi_{\sigma^{\prime}}(\alpha))$
      $\alpha=(z,\hat{c},\hat{y},y^{\prime})$
      $y^{\prime}\sim\text{Categorical}(v),\ v\sim\mathcal{U}\{0,|y|\}$
@@
      Sample to predict counterfactual labels:
      $\hat{y}^{\prime}\sim\text{Categorical}(f(\hat{c}^{\prime})),\ \hat{c}^{\prime}\sim\text{Bernoulli}(\phi_{z}(z^{\prime}))$
Concept-based counterfactuals may also be explained via the interpretable model $f$. The following example illustrates a concrete scenario, also represented in Diagram 2.
Example. A CF-CBM address three questions by design: Predict concept $\hat{c}=\{\hat{c}_{yellow}=1,\hat{c}_{round}=0\}$ and class $\hat{y}=\{\hat{y}_{apple}=0,\hat{y}_{banana}=1\}$ labels. Explain that the class prediction is $\hat{y}_{banana}=1$ because $\hat{c}_{yellow}=1$ and $\hat{c}_{round}=0$ are relevant for the task predictor $f$. Imagine the counterfactual $\{\hat{c}_{yellow}^{\prime}=1,\hat{c}_{round}^{\prime}=1\}$ leading to the different class predictions $\hat{y}^{\prime}=\{\hat{y}_{apple}^{\prime}=1,\hat{y}_{banana}^{\prime}=0\}$.
@@ 3.4 Test-time Functionalities
Counterfactuals explain the effect of concept interventions on downstream tasks.
Plain concept-based explanations indicating the presence or absence of a concept for a given class prediction. However, the complexity of plain explanations grows quickly with the number of concepts. Concept-based counterfactuals, instead, induce simpler, sparser explanations (via Eq. 7) representing minimal modifications of concept labels that would have led to a different class prediction (Abid et al., 2022).
Act upon concept-based counterfactuals.
Concept-based counterfactuals can guide users towards achieving desired outcomes (indicated by class labels), especially when users do not know how to alter their status (represented by concepts). In this scenario: (i) a user specifies a desired class label $y^{\prime}$ for the downstream task representing their goal (e.g., get a loan), (ii) CF-CBMs generate a concept-based counterfactual $p(c^{\prime}|y^{\prime},\hat{y},\hat{c})$ representing minimal concept modifications changing the class label from the predicted $\hat{y}$ to the desired $y^{\prime}$ (e.g., save $10\%$ more each month), (iii) the user can act upon the counterfactual $c^{\prime}$ to accomplish the goal represented by $y^{\prime}$.
Task-driven interventions fix mispredicted concepts.
A key feature of CBMs is enabling human-in-the-loop interventions: at test time, users can correct mispredicted concept labels to enhance downstream task performance ($c\rightarrow y$ intervention). Concept interventions can be useful when intervening on concepts is easier than intervening on the downstream task. For instance, it might be easier to identify bone spurs ($c$) in a knee X-ray scan than to provide an accurate arthritis level ($y$). While still supporting concept interventions, CF-CBMs may also invert this mechanism via task-driven interventions ($y\rightarrow c$), exploiting its counterfactual generation abilities. Task-driven interventions can be used to correct mispredicted class labels when intervening on the downstream task is easier than intervening on concepts. For instance, it might be easier to identify a mispredicted Parkinson’s disease ($y$) rather than intervening in the genetic pathways ($c$) leading to the disease. In this scenario, users can intervene on the class label (rather than on the concepts) suggesting a correct $y^{\prime}$. Considering this additional information, CF-CBMs propose a more accurate set of concept labels $\hat{c}^{\prime}$ representing potential concept intervention previously unknown to the user.
Take the best bet or contemplate a multiverse?
@@ 4 Experiments
  • Why? explanation quality: Can CF-CBMs produce simple and robust explanations?
@@
  • What if? counterfactual actionability: Can CF-CBMs produce valid, plausible, efficient counterfactuals? Can CF-CBMs generate accurate task-driven interventions?
@@
This section describes essential information about experiments. We provide further details in Appendix B.
4.1 Data & task setup
@@
In our experiments we use three different datasets commonly used to evaluate CBMs: dSprites (Matthey et al., 2017), where the task is to predict specific combinations of objects having different shapes, positions, sizes and colours; MNIST addition (Manhaeve et al., 2018), where the task is to predict the sum of two digits; and CUB (Welinder et al., 2010), where the task is to predict bird species based on bird characteristics. In all experiments, images are encoded using a pre-trained ResNet18 (He et al., 2015).
4.2 Evaluation
In our analysis we use the following metrics. What? (prediction generalization): we compute the Area Under the Receiver Operating Characteristic Curve (Hand and Till, 2001) for concepts and tasks (ROC AUC ($\uparrow$)111($\uparrow$): the higher the better. ($\downarrow$): the lower the better.). Why? (explanation quality): we compute the impact of each concept on the downstream task, inspecting the task predictor $f$ (Yuksekgonul et al., 2022), and the impact of intervening on confounding concepts for class predictions (Koh et al., 2020). What if? (counterfactual actionability): drawing from previous works on counterfactuals, we compute: (i) the validity [^2] A.k.a. “correctness” (Abid et al., 2022).($\uparrow$(Wachter et al., 2017) by checking whether the model predicts the desired class labels based on the generated counterfactual; (ii) the proximity ($\downarrow$(Pawelczyk et al., 2020) which evaluates counterfactuals’ “reliability” as their similarity w.r.t. training samples; (iii) the time ($\downarrow$)  (Romashov et al., 2022) to generate a counterfactual for a set of samples.
@@ 4.2 Evaluation
In addition, we propose the following metrics to assess the quality of counterfactuals: (i) the $\Delta$-Sparsity ($\downarrow$) which evaluates user’s “wasted efforts” by counting the number concepts changed (Guo et al., 2023) w.r.t. the minimal number of changes that would have generated a counterfactual according to the dataset; (ii) the “plausibility” (Wachter et al., 2017) as the Intersection over Union (IoU ($\uparrow$)) (Jaccard, 1912) between counterfactuals and ground-truth concept vectors; (iii) the variability ($\uparrow$) which evaluates counterfactuals’ “diversity”, as the cardinality ratio between the set of counterfactuals generated and the set of training concept vectors; (iv) finally, we measure the concept accuracy of generated interventions (Acc Int. ($\uparrow$)) w.r.t. ground-truth optimal interventions that evaluate the model’s ability to generate potential concept interventions. Following Espinosa Zarlenga et al. (2022), we inject noise on predicted concepts to reduce task accuracy, then we sample counterfactuals, conditioning on the ground-truth label, and use them to fix mispredicted concept labels automatically.
All metrics are reported using the mean and the standard error over five different runs with different initializations.
4.3 Baselines
@@
In our experiments, we compare CF-CBMs with standard CBMs in terms of generalization performance. To ensure a fair comparison, we design counterfactual baselines by adapting post-hoc methods to standard CBMs. We propose this experimental design for two main reasons. First, standard CBMs cannot generate counterfactuals by design, hence they need an external generator. Second, the standard, direct application of post-hoc methods on unstructured data types (e.g., image pixels rather than concepts) would have significantly compromised counterfactuals’ interpretability. We remark that this already represents an undocumented procedure in the current literature, as post-hoc methods typically derive counterfactuals from input features, not in the concept space. Among post-hoc approaches we select: (i) the Bayesian Counterfactual (BayCon) (Romashov et al., 2022), which generates counterfactuals via probabilistic feature sampling and Bayesian optimization; (ii) the Counterfactual Conditional Heterogeneous Variational AutoEncoder (C-CHVAE) (Pawelczyk et al., 2020), which iteratively perturbs the latent space of a VAE until it finds counterfactuals; (iii) the VAE CounterFactual (VAE-CF) (Mahajan et al., 2020), which improves C-CHVAE by conditioning counterfactuals on model’s predictions; (iv) the Variational Counter Net (VCNet) (Guyomard et al., 2023), which conditions latent counterfactual sampling on a specific target label $y^{\prime}$.
@@ 5.1 What? Prediction Generalization
CF-CBMs match generalization performance of standard CBMs.
@@
CF-CBMs attain both concept and task ROC AUC similar to standard CBMs (details about these results in Appendix C). This finding demonstrates that generating counterfactuals does not have a negative impact on classification performances on the datasets we considered. This outcome highlights how CF-CBMs generalize standard CBM architectures by matching their predictive accuracy while introducing new fundamental functionalities (i.e., counterfactual explanations and task-driven interventions).
5.2 Why? Explanation Quality
CF-CBMs provide simple and robust concept-based explanations for their predictions (Figure 3).
Following Koh et al. (2020), CF-CBMs provide concept-based explanations showing how each concept contributes to the downstream task prediction (Figure 3). To assess the impact of CF-CBMs design on the explanation’s quality we experiment using a modified version of the dSprites dataset, following Goyal et al. (2019). First, we introduced a high correlation between an object’s color and its class label. Then we removed this correlation on test images, and we checked whether the objects’ color was relevant for class predictions and whether intervening on objects’ colors impacted class predictions. We observed that standard CBM explanations considered both shapes and colors equally important for the class prediction, while CF-CBMs correctly focused on shapes only. Additionally, interventions on an object’s colors influenced $\sim 15\%$ of class predictions in standard CBMs, showing how the model was relying on confounders to solve the task via learning shortcuts (Marconato et al., 2023). CF-CBMs’ predictions, instead, were not affected by interventions on an object’s color, demonstrating better robustness to confounding factors. This robustness stems from CF-CBM’s design to generate a counterfactual altering the minimal number of concept labels to change the class label (e.g., changing the shape but not the color), leading the task predictor $f$ to prioritize the minimal set of decisive features. Appendix C provides further details.
(a) CBM
(b) CF-CBM
Figure 3: CF-CBM produce simpler and more robust concept-based explanations. The plots show two concept-based explanations representing the concept relevance given by the task predictor $f$ for the downstream task on the dSprites data set. Here, class labels depend on objects’ shapes only. Standard CBMs almost equally rely on both objects’ shapes and colors, while CF-CBMs mostly rely on shapes only.
5.3 What if? Counterfactual actionability
dSprites ($\uparrow$) MNIST add ($\uparrow$) CUB ($\uparrow$) avg. ($\uparrow$)
CBM+BayCon $\mathbf{100.0\pm 0.0}$ $94.6\pm 0.5$ $66.2\pm 0.5$ $86.9\pm 0.3$
CBM+CCHVAE $98.4\pm 0.6$ $94.7\pm 0.2$ $78.7\pm 0.4$ $90.6\pm 1.2$
CBM+VAECF $\mathbf{100.0\pm 0.0}$ $82.0\pm 5.9$ $95.2\pm 0.2$ $92.4\pm 2.0$
CBM+VCNET $\mathbf{100.0\pm 0.0}$ $89.9\pm 0.9$ $93.1\pm 0.5$ $94.3\pm 0.5$
CF-CBM (ours) $\mathbf{100.0\pm 0.0}$ $\mathbf{96.4\pm 0.7}$ $\mathbf{99.0\pm 0.1}$ $\mathbf{98.5\pm 0.3}$
@@ 5.3 What if? Counterfactual actionability
CF-CBMs outperform baselines in counterfactuals’ validity. (Table 1)
CF-CBMs attain the highest counterfactual validity across all datasets when compared to counterfactual generation baselines applied to the concept space of standard CBMs. This result shows that CF-CBMs can effectively generate concept vectors that lead to user-requested class labels. The gap w.r.t. baselines increases in complex datasets involving a vast search space over multiple concepts and classes. In the MNIST addition dataset (20 concepts, 19 classes), CF-CBMs achieve $+14$ percentage points (p.p.) higher validity than VAECF and $+7$ p.p. higher than VCNet. In CUB (118 concepts, 200 classes), CF-CBMs find solutions with a significantly higher validity w.r.t. iterative methods: $+33$ p.p. higher than BayCon and $+20$ p.p. higher than C-CHVAE.
Figure 4: CF-CBMs balance the trade off between counterfactuals’ reliability (proximity) and user effort ($\Delta$-Sparsity). The arrow ()normal-→($\rightarrow$)( → ) points towards optimal values. Pareto-optimal models form the frontier of the shaded region, whereas dominated solutions are located within the shaded region.
Figure 5: CF-CBMs calibrate the trade off between counterfactuals’ diversity (variability) and plausibility (IoU). The arrow ()normal-→($\rightarrow$)( → ) points towards optimal values. Pareto-optimal models form the frontier of the shaded region, whereas dominated solutions are located within the shaded region. “CF-CBM $n$” samples $n$ counterfactuals.
Figure 6: CF-CBMs generate reliable (proximity) and accurate interventions (ROC AUC Int.). The arrow ()normal-→($\rightarrow$)( → ) points towards optimal values. Pareto-optimal models form the frontier of the shaded region, whereas dominated solutions are located within the shaded region.
CF-CBMs balance counterfactuals’ reliability with minimal user effort. (Figure 4)
High-quality counterfactuals should satisfy two key properties simultaneously: to be reliable and to require the least amount of users’ effort to act upon. However, minimizing user’s effort alone might lead to learning shortcuts producing less reliable counterfactuals, while optimizing reliability might generate counterfactuals that change the values of multiple concepts and represent a significant burden for users to take actions. For this reason, these two properties should be evaluated together:

### v3

@@ 3. Why not? Imagine interpretable counterfactuals via the generation of new concept tuples:
    $z^{\prime}\sim\mathcal{N}(\phi_{\mu^{\prime}}(\alpha),\phi_{\sigma^{\prime}}(\alpha)),\quad\alpha=(z,\hat{c},\hat{y},y^{\prime}),\quad y^{\prime}\sim\text{Categorical}(v),\ v\sim\mathcal{U}\{0,|y|\}$
@@
    Sample to predict counterfactual labels: $\hat{y}^{\prime}\sim\text{Cat}(f(\hat{c}^{\prime})),\ \hat{c}^{\prime}\sim\text{Ber}(\phi_{z}(z^{\prime}))$
The following example illustrates a concrete scenario.
[…]
@@ 3.3 Test-time Functionalities
Counterfactuals explain the effect of concept interventions on downstream tasks. Plain concept-based explanations indicate the presence or absence of a concept for a given class prediction. However, the complexity of plain explanations grows quickly with the number of concepts. Concept-based counterfactuals, instead, induce simpler, sparser explanations (via Eq. 5) representing minimal modifications of concept labels that would have led to a different class prediction.
Act upon concept-based counterfactuals. Concept-based counterfactuals can guide users towards achieving desired outcomes (indicated by class labels), especially when users do not know how to alter their status (represented by concepts). In this scenario: (i) a user specifies a desired class label $y^{\prime}$ for the downstream task representing their goal (e.g., get a loan), (ii) CF-CBMs generate a concept-based counterfactual $p(c^{\prime}|y^{\prime},\hat{y},\hat{c})$ representing minimal concept modifications changing the class label from the predicted $\hat{y}$ to the desired $y^{\prime}$ (e.g., save $10\%$ more each month), (iii) the user can act upon the counterfactual $c^{\prime}$ to accomplish the goal represented by $y^{\prime}$.
Task-driven interventions fix mispredicted concepts. A key feature of CBMs is enabling human-in-the-loop interventions: at test time, users can correct mispredicted concept labels to enhance downstream task performance ($c\rightarrow y$ intervention). Concept interventions can be useful when intervening on concepts is easier than intervening on the downstream task. For instance, it might be easier to identify if the lung is collapsed ($c$) in a lung X-ray scan than to provide an accurate prediction if a patient has or not pneumothorax ($y$). While still supporting concept interventions, CF-CBMs may also invert this mechanism via task-driven interventions ($y\rightarrow c$), exploiting its counterfactual generation abilities. Task-driven interventions can be used to correct mispredicted class labels when intervening on the downstream task is easier than intervening on concepts. For instance, it might be easier to identify a mispredicted Parkinson’s disease ($y$) rather than intervening in the genetic pathways ($c$) leading to the disease. In this scenario, users can intervene on the class label (rather than on the concepts) suggesting a correct $y^{\prime}$. Considering this additional information, CF-CBMs propose a more accurate set of concept labels $\hat{c}^{\prime}$ representing potential concept intervention previously unknown to the user.
@@ 4 Experiments
  • How? intervention impact: Does optimizing counterfactual generation end-to-end with the predictor alter the importance assigned by the model to each concept?
@@
  • Why not? counterfactual actionability: Can CF-CBMs produce valid, plausible, efficient counterfactuals? Can CF-CBMs generate accurate task-driven interventions?
@@
This section describes essential information about experiments. We provide further details in Appendix C.
Data & task setup
@@
In our experiments we use five different datasets commonly used to evaluate CBMs: dSprites (Matthey et al., 2017), where the task is to predict specific combinations of objects having different shapes, positions, sizes and colours; MNIST addition (Manhaeve et al., 2018), where the task is to predict the sum of two digits; and CUB (Welinder et al., 2010), where the task is to predict bird species based on bird characteristics; CIFAR10 (Krizhevsky et al., ), where the task is to classify the object in the image, and SIIM Pneumothorax (Zawacki et al., 2019), where the task is to determine whether the X-ray scan indicates Pneumothorax. These two datasets do not include concept annotations, so we extract them following the method proposed in Oikarinen et al. (2023). In all experiments, images are encoded using a pre-trained ResNet18 (He et al., 2015) (dSpties, MNIST Add, CUB), CLIP ViTB16 (Radford et al., 2021)(CIFAR10) or CXR-CLIP (You et al., 2023) (SIIM Pneumothorax).
Evaluation
@@
In our analysis we use the following metrics. What? (prediction generalization): we compute the Area Under the Receiver Operating Characteristic Curve (Hand & Till, 2001) for concepts and tasks (ROC AUC ($\uparrow$)222($\uparrow$): the higher the better. ($\downarrow$): the lower the better.). How? (intervention impact): we assess the influence of concepts on the downstream task using the Causal Concept Effect (CaCE) (Goyal et al., 2019), which quantifies the degree to which each concept is critical for a class. This allows us to evaluate the impact of optimizing counterfactuals end-to-end versus using post-hoc methods. Additionally, in scenarios where confounders are present, CaCE can help determine whether the impact of these confounders is high or low. Why not? (counterfactual actionability): drawing from previous works on counterfactuals, we compute: (i) the validity [^3] A.k.a. “correctness” (Abid et al., 2022).($\uparrow$(Wachter et al., 2017) by checking whether the model predicts the desired class labels based on the generated counterfactual; (ii) the proximity ($\downarrow$(Pawelczyk et al., 2020) which evaluates counterfactuals’ “reliability” as their similarity w.r.t. training samples; (iii) the time ($\downarrow$)  (Romashov et al., 2022) to generate counterfactuals (see Appendix D for the results on this metric). In addition, we propose the following metrics to assess the quality of counterfactuals: (i) the $\Delta$-Sparsity ($\downarrow$) which evaluates user’s “wasted efforts” by counting the number concepts changed (Guo et al., 2023) w.r.t. the minimal number of changes that would have generated a counterfactual according to the dataset; (ii) the “plausibility” (Wachter et al., 2017) as the Intersection over Union (IoU ($\uparrow$)) (Jaccard, 1912) between counterfactuals and ground-truth concept vectors; (iii) the variability ($\uparrow$) which evaluates counterfactuals’ “diversity”, as the cardinality ratio between the set of counterfactuals generated and the set of training concept vectors; (iv) finally, we measure the concept accuracy of generated interventions (Acc Int. ($\uparrow$)) w.r.t. ground-truth optimal interventions that evaluate the model’s ability to generate potential concept interventions. Following Espinosa Zarlenga et al. (2022), we inject noise on predicted concepts to reduce task accuracy, then we sample counterfactuals, conditioning on the ground-truth label, and use them to fix mispredicted concept labels automatically. All metrics are reported using the mean and the standard error over five different runs with different initializations. Additionally, in Appendix E, we provide examples of counterfactuals generated by our models and the baselines across various datasets.
Baselines
@@
In our experiments, we compare CF-CBM with a Black Box model, a standard CBM and a more powerful CBM as Concept Embedding Model (CEM) (Espinosa Zarlenga et al., 2022) in terms of generalization performance. To ensure a fair comparison, we create counterfactual baselines by adapting post-hoc methods to standard CBMs, allowing for a direct comparison between our approach and other state-of-the-art models. We propose this experimental design for two main reasons. First, standard CBMs cannot generate counterfactuals by design, hence they need an external generator. Second, the standard, direct application of post-hoc methods on unstructured data types (e.g., image pixels rather than concepts) would have significantly compromised counterfactuals’ interpretability and difficult comparison between the generated results. We remark that this already represents an undocumented procedure in the current literature, as post-hoc methods typically derive counterfactuals from input features, not in the concept space. However, this also enables us to compare post-hoc counterfactual methods with our model, which incorporates counterfactual constraints during training. Among post-hoc approaches we select: (i) the Bayesian Counterfactual (BayCon) (Romashov et al., 2022), which generates counterfactuals via probabilistic feature sampling and Bayesian optimization; (ii) the Counterfactual Conditional Heterogeneous Variational AutoEncoder (C-CHVAE) (Pawelczyk et al., 2020), which iteratively perturbs the latent space of a VAE until it finds counterfactuals; (iii) the VAE CounterFactual (VAE-CF) (Mahajan et al., 2020), which improves C-CHVAE by conditioning counterfactuals on model’s predictions; (iv) the Variational Counter Net (VCNet) (Guyomard et al., 2023), which conditions latent counterfactual sampling on a specific target label $y^{\prime}$.
@@ 5.1 What? Prediction Generalization
CF-CBMs achieve generalization performance that is close to standard CBMs. CF-CBMs attain both concept and task ROC AUC similar to Black Box, standard CBM and CEM, a more expressive CBM (details in Appendix D). This shows that generating counterfactuals does not have a negative impact on classification performance on the datasets we considered. Our goal was not to improve performance but to improve interpretability. This outcome highlights how CF-CBMs generalise standard CBM architectures by matching their predictive accuracy while introducing new fundamental functionalities (i.e., counterfactual explanations and task-driven interventions).
5.2 How? Intervention impact
Table 1: CF-CBMs ignore the confounding concept on the modified dSprites dataset.
Post-hoc Joint (ours)
Shape 1 $0.16\pm 0.03$ $0.30\pm 0.15$
Shape 2 $0.13\pm 0.03$ $0.18\pm 0.08$
Shape 3 $0.16\pm 0.03$ $0.32\pm 0.09$
Two obj $0.04\pm 0.01$ $0.16\pm 0.07$
Colours (confounder) $0.17\pm 0.03$ $0.02\pm 0.01$
Table 2: Compared to CBMs, CF-CBMs assign greater importance to a small fraction of concepts, while the majority of concepts receive less importance on CUB.
Dataset Percentile Mean Post-hoc Mean Joint (ours) p-value
MNIST Top 10% ($\uparrow$) $0.1589\pm 0.0203$ $\textbf{0.1827}\pm 0.0676$ 0.0001
Bottom 50% ($\downarrow$) $0.0308\pm 0.0175$ $\textbf{0.0257}\pm 0.0143$ $0.2437$
CUB Top 10% ($\uparrow$) $0.0093\pm 0.0029$ $\textbf{0.0131}\pm 0.0044$ 0.0
Bottom 50% ($\downarrow$) $0.0016\pm 0.0006$ $\textbf{0.0013}\pm 0.0008$ 0.0
Jointly training counterfactuals’ generator and CBMs makes the model rely on fewer important concepts leading to simpler explanations (Table 2) Following Koh et al. (2020), CF-CBMs allow users to intervene directly on the concepts, enabling simulation of diverse scenarios and observation of their effects on the model’s final prediction, as quantified by the Causal Concept Effect (CaCE) (Goyal et al., 2019). To assess the impact of CF-CBMs joint training design on how the model perceives the importance of each concept we conducted experiments using a modified dSprites dataset, following Goyal et al. (2019), where a high correlation was introduced between an object’s color and its class label. While standard CBM explanations weighted shape and color (confounders) equally for class prediction, CF-CBMs correctly prioritized shape as the primary predictor, as shown in Table 2. Specifically, altering color in standard CBMs resulted in a CaCE score of 0.17, demonstrating a 17% chance of prediction change due to this confounding factor. This reveals that standard CBMs are susceptible to relying on shortcuts in decision-making (Marconato et al., 2023). In contrast, CF-CBMs showed minimal sensitivity to color changes, with a CaCE score of just 0.02, demonstrating significantly higher robustness against confounding features. Additionally, CF-CBMs assigned higher CaCE scores to shape concepts, indicating a stronger reliance on these critical features, unlike the post-hoc version. This trend was consistently observed across the other datasets: MNIST (p-value = 0.003) and CUB (p-value = 0.000), where CF-CBMs and CBMs exhibited distinct CaCE distributions as confirmed by the Kolmogorov-Smirnov test (the distributions are shown in App. D). Table 2 further illustrates that CF-CBMs achieved a higher mean CaCE for the top 10% of concepts, signifying that the model concentrates on a smaller subset of highly relevant concepts. This focus results in more concise and interpretable explanations by engaging a reduced number of concepts. Moreover, CF-CBMs exhibited a lower mean CaCE for the bottom 50% of concepts, suggesting that the model learned to disregard less critical features, thereby enhancing robustness against concept perturbations. The Mann-Whitney U test results reveal that three out of four comparisons were statistically significant, underscoring the impact of joint training. This shift in decision-making comes from the CF-CBM’s objective of generating counterfactuals that alter only the minimal subset of concept labels required to change the class prediction—e.g., adjusting shape but not color. Consequently, the model’s task predictor is forced to focus on the minimal set of decisive features. For additional details on the CaCE score distributions, refer to Appendix D. Overall, our findings suggest that jointly training CF-CBMs not only improves model interpretability but also yields a more robust decision-making process.
@@ 5.2 How? Intervention impact
dSprites ($\uparrow$) MNIST add ($\uparrow$) CUB ($\uparrow$) CIFAR10 ($\uparrow$) SIIM Pneumothorax ($\uparrow$) avg. ($\uparrow$)
CBM+BayCon $\mathbf{100.0\pm 0.0}$ $94.6\pm 0.5$ $66.2\pm 0.5$ $91.0\pm 3.3$ $\mathbf{100.0\pm 0.0}$ $90.4\pm 0.7$
CBM+CCHVAE $98.4\pm 0.6$ $94.7\pm 0.2$ $78.7\pm 0.4$ $93.6\pm 0.1$ $93.0\pm 0.6$ $91.7\pm 0.2$
CBM+VAECF $\mathbf{100.0\pm 0.0}$ $82.0\pm 5.9$ $95.2\pm 0.2$ $91.0\pm 0.8$ $\mathbf{100.0\pm 0.0}$ $93.4\pm 1.2$
CBM+VCNET $\mathbf{100.0\pm 0.0}$ $89.9\pm 0.9$ $93.1\pm 0.5$ $\mathbf{100.0\pm 0.0}$ $\mathbf{100.0\pm 0.0}$ $96.6\pm 0.2$
CF-CBM (ours) $\mathbf{100.0\pm 0.0}$ $\mathbf{96.4\pm 0.7}$ $\mathbf{99.0\pm 0.1}$ $\mathbf{100.0\pm 0.0}$ $\mathbf{99.9\pm 0.1}$ $\mathbf{99.06\pm 0.1}$
5.3 Why not? Counterfactual actionability
CF-CBMs outperform baselines in counterfactuals’ validity. (Table 3) CF-CBMs attain the highest counterfactual validity across all datasets when compared to counterfactual generation baselines applied to the concept space of standard CBMs. This result shows that CF-CBMs can effectively generate concept vectors that lead to user-requested class labels. The gap w.r.t. baselines increases in complex datasets involving a vast search space over multiple concepts and classes. In the MNIST addition dataset (20 concepts, 19 classes), CF-CBMs achieve $+14$ percentage points (p.p.) higher validity than VAECF and $+7$ p.p. higher than VCNet. In CUB (118 concepts, 200 classes), CF-CBMs find solutions with a significantly higher validity w.r.t. iterative methods: $+33$ p.p. higher than BayCon and $+20$ p.p. higher than C-CHVAE. Lastly, in the CIFAR10 and SIIM Pneumothorax datasets, CF-CBMs continue to achieve the highest validity.
CF-CBMs balance counterfactuals’ reliability with minimal user effort. (Figure 3) High-quality counterfactuals should satisfy two key properties simultaneously: to be reliable and to require the least amount of users’ effort to act upon. However, minimizing user’s effort alone might lead to learning shortcuts producing less reliable counterfactuals, while optimizing reliability might generate counterfactuals that change the values of multiple concepts and represent a significant burden for users to take actions. For this reason, these two properties should be evaluated together: